OpenStudy (anonymous):

17 + 2 [8 - (-3)^3] is 87?

OpenStudy (angela210793):

17+2*(8+27)=17+70=87 :) Gj :D

OpenStudy (saifoo.khan):

Righto^

OpenStudy (angela210793):

wht happened to ''Agree^^''?

OpenStudy (saifoo.khan):

Agreed as well.

OpenStudy (saifoo.khan):

^_^

OpenStudy (angela210793):

^_^

OpenStudy (anonymous):

i messed up the first time i wrote the problem up so had to redo it